OpenStudy (anonymous):

Which of the following statements best describes the graph of 3x – 2y = 4? It is a straight line joining the points (6, 2), (5, 1), and (7, 3). It is a straight line joining the points (0, –2), (2, 1), and (–2, –5). It is a curve joining the points (3, –2), (2, 3), and (4, 1). It is a curve joining the points (0, 2), (–2, –1), and (4, 1).

OpenStudy (anonymous):

Okay, the equation can be written in the from y = mx + c where m is the gradient and c is a constant of any value. This means that it's a straight line. To see whether it is a or b, simply input the x and y values they give you into the equation: 3(4) - 2(2) = 4 12 - 4 = 4 8 = 4 which is incorrect so it can't be the first one meaning it must be the second.

OpenStudy (anonymous):

Thanks alot